OpenStudy (anonymous):

I need help with this ACT review question. I already finished this and didn't answer this question. How does one solve this?

OpenStudy (anonymous):

Here

geerky42 (geerky42):

Say we have sin x. What's the value x have to be for sin x to reach its maximum (1)?

OpenStudy (anonymous):

I have no idea, Geometry was a couple years ago lol. So far I've noticed this type of trig is raping me

geerky42 (geerky42):



geerky42 (geerky42):

Do you remember something about that? sinx is basically height of triangle in unit circle So what's degree x have to be for this triangle to reach highest possible height?

OpenStudy (anonymous):

90?

OpenStudy (anonymous):

All I'm thinking in my head is Soh Cah Toa and it doesn't seem to do anything for this lol, sorry I'm so bewildered

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@ParthKohli @phi @jim_thompson5910 @mathstudent55 @jdoe0001 @zepdrix @campbell_st

zepdrix (zepdrix):

sup? :U

OpenStudy (anonymous):

Lol, you up to help me?

zepdrix (zepdrix):

So uhhh\[\Large\rm y=\sin2a\]We want the a values that give us \(\Large\rm y=1\) because that's the maximum for the sine function. And we want the smallest a value which does so. Where we gettin stuck? Evaluating specific a's? So for a=pi/4, \[\Large\rm y=\sin2 \frac{\pi}{4}\]\[\Large\rm y=\sin\frac{\pi}{2}\]Do you know sin(pi/2)? :O

OpenStudy (anonymous):

Alright I did not remember that y=1 part, let's see

OpenStudy (anonymous):

Isn't that 1.57...?

zepdrix (zepdrix):

Sin(pi/2) ? No. Again, remember that the `maximum` of the sine function is 1. So you should be getting something like 1, or a decimal less than 1.

zepdrix (zepdrix):

If you're putting it into the calculator, make sure you're in `radians` mode.

OpenStudy (anonymous):

Oh ok, I wasn't even sure what I should have been doing. I put it in the calculator and got 1

zepdrix (zepdrix):

Ok great. So a=pi/4 ended up giving us y=1, yes?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

We're looking for the `smallest` a value that will give us y=1. So we should check the other ones if they're smaller than pi/4. Do we have any others that we need to check? Any other a values smaller than pi/4?

OpenStudy (anonymous):

No sir

zepdrix (zepdrix):

Mmm ok great! So it seems a=pi/4 is the correct choice! That is the smallest a value (among our options) that will maximize the sine function.

OpenStudy (anonymous):

Alright so y has to equal 1 for sin, is there something else for the other ones?

zepdrix (zepdrix):

y has to equal 1 in order to be a `maximum of the sine function`. Other what? >.<

OpenStudy (anonymous):

Cos, etc?

zepdrix (zepdrix):

Cosine also has a maximum of 1. The other functions beyond that are a little weird though :o